OpenStudy (anonymous):

A helicopter is climbing at 8.0m/s when it drops a pump near a leaking boat. The pump reaches the water 4.0s afterward. How high was the helicopter when the pump was dropped? When the pump reached the water? What formula should I use?

OpenStudy (anonymous):

S=ut+gt^2/2, where u is initial speed, g is acceleration of gravity and t is time. Since the pump fell starting from rest, then s=gt^2/2 or 9.81*4^2/2=78.48metres.

OpenStudy (anonymous):

Aren't you gonna plug in 8 m/s to u?

OpenStudy (anonymous):

No, do not think so, because that is just the speed of the helicopter itself. I bet that was the trick. It is like you bungee jumping and while you are in air and spit towards the ground, your saliva will have an initial speed of 0m/s, not the one you were speeding in.

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

No problem.